white chocolate powder worth 4 pesos and 10 cents per pound is to be mixed with the black chocolate powder worth 4 pesos and 90 cents per pound.How many pound of white and black chocolate should used to obtain 25 pounds of blended chocolate worth 4 pesos and 40 cents per pound.

Answer by stanbon(75887)

white chocolate powder worth 4 pesos and 10 cents per pound is to be mixed with the black chocolate powder worth 4 pesos and 90 cents per pound.How many pound of white and black chocolate should used to obtain 25 pounds of blended chocolate worth 4 pesos and 40 cents per pound.
------
Weight Eq:: w + c = 25 lbs
Value Eq:: 4.1w + 4.9c = 4.4*25
--------------------
41w + 41c = 41*25
41w + 49c = 44*25
---------
Subtract and solve for "c":
8c = 3*25
c = 75/8 = 9.375 lb (amt. of black chocolate to use)
----
w = 25 - 9.375 = 15.625 lb (amt. of white chocolate to use)
-------------
Cheers,
Stan H.
-------------